Starting population of mice N = 1000 During the 1st month B = 200 D = 50 Q1. What is b, for the first month?

Starting population of mice, b & d values for 1 month N = 1000 b = 0.1 d = 0.2 Q2. What is N after 1 month?
Answer #1
  1. Net population(N) = 1000, Birth(B) = 200 and Death(D) = 50.............during 1st month

% of the population were birth (b) :

(Birth/Net population) *100 = (200/1000)*100 = 0.2 or 20%

2. Given Net population(N) = 1000, b= 0.1, d = 0.2

Net population after 1 month;

Net population * b = 1000 * 0.1 = 100

It means 100 new individual is added after 1 month.